# supplementary

### v1

@@ 1 Introduction
Recent advances in visual-language models (VLMs), as exemplified by CLIP (Radford et al. 2021), ALIGN (Jia et al. 2021), BLIP (Li et al. 2022b), Flamingo (Alayrac et al. 2022), LLaVA (Liu et al. 2023), and MiniGPT-4 (Zhu et al. 2024), have significantly narrowed the long-standing gap between visual perception and language understanding. By leveraging large-scale image-text pairs during pre-training, those VLMs exhibit remarkable capabilities on various real-world visual understanding tasks such as classification (Radford et al. 2021; Yang et al. 2022), object detection (Li et al. 2022c; Zhong et al. 2022), segmentation (Li et al. 2022a; Zhang et al. 2023), captioning (Li et al. 2023a), visual question answering (Alayrac et al. 2022; Li et al. 2022b; Bai et al. 2023; Wang et al. 2024), and visual reasoning (Liu et al. 2023).
@@
Prompt tuning (Zhou et al. 2022b, a; Tian et al. 2024; Li et al. 2024) has emerged as a parameter-efficient approach to fine-tune VLMs. The introduction of a few learnable prompts to a frozen pre-trained VLM allows it to adapt to various downstream tasks. However, those prompts tend to overfit the task-specific data distributions, raising concerns about their generalization capabilities across diverse tasks.
@@
To address the limited generalization capability of prompts, we consider the label smoothing (LS) (Szegedy et al. 2016; Müller, Kornblith, and Hinton 2019; Yuan et al. 2020), a widely recognized regularization technique. LS was originally introduced to mitigate overfitting in classification tasks. It improves generalization by distributing a small portion of probability mass across incorrect labels, which discourages model over-confidence and encourages softer, more calibrated predictions (Zhang et al. 2021; Liang et al. 2022; Lukasik et al. 2020). Motivated by its effectiveness in reducing over-fitting and enhancing generalization, we explore the integration of LS with prompt tuning.
@@
Figure 1: Comparison of test accuracies among CoOp (Zhou et al. 2022b), CoOp with vanilla label smoothing (LS) (Szegedy et al. 2016), TFKD (Yuan et al. 2020), OLS (Zhang et al. 2021), and alternating training-based label smoothing (ATLaS), where prompts are trained on the base classes of ImageNet and subsequently evaluated on new classes.
@@
However, our preliminary investigation reveals that employing the LS strategy for prompt tuning fails to enhance the generalization capability of prompts, which aligns with an observation in (Kobyzev et al. 2022) that “label smoothing doesn’t improve fine-tuning of a pre-trained model”. Specifically, we conduct preliminary experiments using a representative prompt tuning method, CoOp (Zhou et al. 2022b), as a baseline, where experimental settings are described in Table 3 of the base-to-new generalization part. As illustrated in Figure 1, CoOp with vanilla LS demonstrates slightly inferior testing performance compared to the CoOp method throughout the prompt tuning process. The variants of LS do not improve the testing performance (i.e., TFKD (Yuan et al. 2020)), but rather impair it (i.e., OLS (Zhang et al. 2021)). Naturally, we raise a question: “Can the LS be revised to help improve the generalization capacity of prompt tuning”.
@@
To answer this question, we propose the Alternating Training-based Label Smoothing (ATLaS) method, a novel label smoothing technique, for prompt tuning. Specifically, ATLaS alternates between using one-hot labels and soft labels as the supervision information to guide prompt tuning, thereby enhancing the generalization capability of learnable prompts. Figure 1 shows that CoOp with ATLaS successfully enhances the prompt generalization. For the generation of soft labels in VLMs, we propose two efficient methods. The first method produces Class-wise Soft Labels (CSL) to derive inter-class similarities from textual prompts, and the second method generates Instance-wise Soft Labels (ISL) to capture instance-class associations from the interaction between textual prompts and instances. Those two types of soft labels could effectively guide the learning of prompts by providing rich supervision signals from different perspectives. Theoretically, we analyze ATLaS from the optimization perspective, demonstrating its advantages over the vanilla LS. Extensive experiments demonstrate that the proposed ATLaS method effectively enhances the generalization of prompts in several settings (i.e., cross-dataset generalization, domain generalization, base-to-new generalization, and few-shot classification).
@@
Our contributions are four-fold. (i) To the best of our knowledge, this is the first attempt to successfully integrate LS into prompt tuning to improve the generalization capacity. To address the ineffectiveness of vanilla LS in prompt tuning, we propose the ATLaS method to guide prompt tuning by alternating between one-hot labels and soft labels; (ii) We introduce two types of soft labels to guide prompt tuning, including class-wise and instance-wise soft labels, both of which can be efficiently generated before the training process; (iii) The convergence properties of the proposed ATLaS method are analyzed; (iv) Extensive experiments demonstrate the effectiveness of the proposed ATLaS method.
2 Related Work
@@
Vision-Language Models. VLMs (Radford et al. 2021; Jia et al. 2021; Li et al. 2022b; Alayrac et al. 2022; Liu et al. 2023; Zhu et al. 2024) have significantly enhanced cognitive capabilities by integrating visual and textual modalities, excelling at real-world vision understanding tasks such as image classification. Among those models, CLIP (Radford et al. 2021) stands out as a foundational pioneer, establishing the standard practice of aligning visual and textual embeddings within a unified representation space. It has influenced the development of subsequent models. For example, Flamingo (Alayrac et al. 2022), LLaVa (Liu et al. 2023), and BLIP-2 (Li et al. 2023b) are built upon the visual-textual alignment encoder introduced by CLIP. Because of its highly representative nature, we use CLIP as the foundational VLM to explore the novel prompt tuning paradigm with label smoothing.
@@
Prompt Tuning. As a lightweight model adaptation method, prompt tuning has attracted widespread attention in VLMs (Zhu et al. 2023; Yao, Zhang, and Xu 2023; Khattak et al. 2023b; Chen, Fu, and Zhang 2024). Unlike traditional fine-tuning methods that update all parameters of the pre-trained model, prompt tuning optimizes only the prompt representations, thereby reducing the computational costs. In particular, CoOp (Zhou et al. 2022b) achieves efficient adaptation by optimizing a set of learnable prompt vectors, which replace the hand-crafted textual prompts (e.g., “a photo of panda”) at the language branch of CLIP. However, despite its simplicity and effectiveness, CoOp tends to overfit base classes, exhibiting limited generalization ability. To address this limitation, CoCoOp (Zhou et al. 2022a) learns a lightweight meta-net that generates input-conditional vectors for each input image. ProGrad (Zhu et al. 2023) regulates the gradient direction to mitigate conflicts between hand-crafted prompts and learnable prompts. MaPLe (Khattak et al. 2023a) introduces a multi-modal prompt learning method that simultaneously learns hierarchical prompts for both vision and language branches to improve the alignment between the visual and textual representations. To address the base-new trade-off issue, DePT (Zhang et al. 2024) decouples base-specific knowledge and task-shared knowledge in feature channels.
@@
Label Smoothing. Empirical evidence proves that applying LS strategies during training improves the generalization of the model (Szegedy et al. 2016). LS is a regularization strategy that replaces one-hot labels with soft labels, preventing the model from becoming over-confident in its predictions. Müller et al. (Müller, Kornblith, and Hinton 2019) demonstrate that LS encourages samples from the same class to group in tight clusters, which results in a loss of information in the logits regarding the similarities of different classes, consequently impairing knowledge distillation. To generate more reliable soft labels, Online label smoothing (OLS) (Zhang et al. 2021) derives class-wise soft labels online by retaining historical predictions. Yuan et al. (Yuan et al. 2020) explores the intrinsic relationship between LS and knowledge distillation, and proposes a teacher-free knowledge distillation (TFKD) framework, where a student model engages in self-learning or draws insights from manually designed regularization distribution. Different from existing LS methods whose optimization paradigm is to jointly optimize a designed loss of label smoothing along with the standard cross-entropy loss, the proposed ATLaS method alternately optimizes the model with the loss of label smoothing and the standard cross-entropy loss, achieving better performance.
3 Methodology
@@
3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP) (Radford et al. 2021) is a VLM pre-trained on about 400 million image-text pairs, resulting in remarkable zero-shot image recognition capability. CLIP enables zero-shot image classification by measuring the similarity between visual and textual embeddings. Specifically, CLIP consists of a visual encoder and a text encoder. Let $\mathbf{I}$ be the visual embedding extracted by the visual encoder in CLIP for an image $\mathbf{x}$, and $\{\mathbf{t}_{c}\}_{c=1}^{C}$ be a set of textual embeddings generated by the text encoder for textual prompts, where $C$ is the number of classes. Those prompts typically follow the format like “a photo of a [CLS]”, where “[CLS]” denotes a class token which is subsequently replaced by specific class names such as “panda”, “dog”, and “bird”. Given those embeddings, the prediction probability for image classification can be computed as
@@ 3.1 Preliminaries
Context Optimization (CoOp) (Zhou et al. 2022b) is a prompt tuning method designed to replace textual prompts with learnable prompts, enabling VLMs to adapt to downstream tasks through parameter-efficient fine-tuning with a few labeled samples. Specifically, CoOp introduces learnable prompt vectors $\mathbf{v}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{M}\}$ to replace the textual prompt (e.g., “a photo of a”), where each $\mathbf{v}_{i}$ has the same dimension as the word embeddings. The learnable prompt is subsequently constructed by concatenating $\mathbf{v}$ with the class token. Formally, the learnable prompts are formed as $\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{M},[\mathrm{CLS}]\}$. Let $\{\mathbf{t}_{c}^{\mathrm{s}}\}_{c=1}^{C}$ be a set of textual embeddings generated by the text encoder for the learnable prompts. The learnable prompt can be optimized by minimizing the cross-entropy loss as
@@
Label Smoothing (LS) (Szegedy et al. 2016) is a regularization strategy that uses soft labels, generated by the weighted averaging of a uniform distribution and hard labels, as a substitute for hard labels during the training process. This technique is designed to mitigate the model over-confidence, thereby enhancing the generalization capabilities of the model. Specifically, LS converts the one-hot label $\mathbf{y}$ into a soft target
@@
where $\theta\in[0,1]$ is a smoothing parameter, and $\hat{\mathbf{y}}$ is drawn from a distribution $\mathbb{P}_{\hat{\mathbf{y}}}$. Following prior works (Szegedy et al. 2016; Müller, Kornblith, and Hinton 2019; Yuan et al. 2020), we consider a uniform distribution across all $C$ classes for $\hat{\mathbf{y}}$, i.e., $\hat{\mathbf{y}}={\frac{1}{C}}\mathbf{1}_{C}$, where $\mathbf{1}_{C}\in\mathbb{R}^{C}$ is the vector of all ones.
3.2 Alternating Training-based Label Smoothing
The prompts learned in CoOp exhibit limited generalization capabilities (Zhou et al. 2022a). As the LS strategy is specifically proposed to enhance model generalization, this naturally motivates the integration of LS into the prompt tuning framework to enhance the generalization of the prompts.
However, as discussed in the introduction (i.e., Figure 1), directly using vanilla LS fails to improve the generalization performance and even impairs it. One possible reason is that the pre-trained CLIP model is trained with one-hot labels but not soft labels, making LS integrated with prompt tuning inconsistent with such a training process and thereby weakening the generalization capacities.
To alleviate the inconsistency of labels used in the pre-training and fine-tuning phases as discussed above, the proposed ATLaS method illustrated in Figure 2 utilizes both the one-hot labels and soft labels. Specifically, ATLaS trains with the one-hot labels and soft labels alternately in $K$ successive training epochs, where the construction of soft labels will be introduced in Section 3.4. That is, in $K$ successive training epochs, the ATLaS method learns the prompts using one-hot labels $\mathbf{y}$ in the first (K1)(K-1) epochs and then trains with the soft labels $\mathbf{y}^{\mathrm{LS}}$ in the $K$-th epoch. This alternating training strategy continues throughout the training process.
3.3 Analysis
@@
We present the following assumptions about smoothness and bounded variance of stochastic gradients, which are commonly used in the non-convex optimization literature (Ghadimi and Lan 2013; Reddi et al. 2016; Yan et al. 2018).
Assumption 1.
@@
As articulated in Theorem 1, a smaller $\kappa$, which is associated with the quality of $\hat{\mathbf{y}}$ employed in LS, facilitates faster convergence. It is noteworthy that the convergence bound for vanilla LS is $\frac{2F(\mathbf{v}_{0})}{\eta T}+(1-\theta+\theta\kappa)\sigma^{2}$ (Xu et al. 2020), [^2] We provide a detailed convergence analysis of label smoothing by (Xu et al. 2020) in Appendix A.4. which is unequivocally larger than that for ATLaS when $\kappa>1$. This suggests that ATLaS can converge better during the optimization of parameters where the selection of $\hat{\mathbf{y}}$ is suboptimal (i.e., characterized by a large $\kappa$).
@@
3.4 Generation of Offline Soft Labels
@@
Due to inherent factors such as pose, deformation, and lighting conditions, images within the same category can exhibit substantial intra-class variance (Xiao et al. 2015; Lu et al. 2022). Such intra-class variance prevents CSL from adequately capturing visual differences. Moreover, the utilization of knowledge in a single modality to supervise prompt tuning could potentially disrupt the vision-language alignment learned by CLIP. To address those concerns in CSL, we propose the ISL, which utilizes both the textual and visual modalities through the interaction between textual prompts and individual instances to capture instance-class associations. Specifically, we can use the prediction probabilities of CLIP as soft labels, which are instance-dependent, and learn the instance-class relations. However, this approach can not guarantee that the value for the ground-truth class in the soft label be the largest. Therefore, the soft label in ISL is rectified to be a combination of the one-hot label $\mathbf{y}$ and the prediction probabilities as
@@
where $\alpha$ is the rectification coefficient, $\mathbf{p}(\cdot|\mathbf{x})=[p(y_{1}|\mathbf{x}),p(y_{1}|\mathbf{x}),\dots,p(y_{C}|\mathbf{x})]^{\mathrm{T}}$ denotes the prediction probabilities by CLIP, $\mathbb{I}(\cdot)$ is the indicator function, and $\delta=\mathbb{I}(\mathrm{arg}\max\mathbf{p}(\cdot|\mathbf{x})\neq\mathrm{the}\;\mathrm{ground}\mbox{-}\mathrm{truth}\;\mathrm{class})$. In other words, the rectification process defined in Eq. (5) is exclusively implemented when the class corresponding to the maximum value of the soft label is not the ground-truth class.
@@
In summary, the whole algorithm for the proposed ATLaS method is shown in Appendix 1.
@@
3:  $\xi:=\begin{cases}0,\ \mathrm{if}\;e+1\equiv 0\left(\mathrm{mod}\;K\right)\\ 1,\mathrm{otherwise}\end{cases}$
@@
4 Experiments
4.1 Experimental Setup

### v2

# Alternating Training-based Label Smoothing Enhances Prompt Generalization

@@ I Introduction
Recent advances in visual-language models (VLMs), as exemplified by CLIP [1], ALIGN [2], BLIP [3], EVA-CLIP [4], LLaVA [5], and MiniGPT-4 [6], have significantly narrowed the long-standing gap between visual perception and language understanding. By leveraging large-scale image-text pairs during pre-training, those VLMs exhibit remarkable capabilities on various real-world visual understanding tasks such as classification [1, 7], object detection [8, 9], segmentation [10, 11], captioning [12], visual question answering [13, 3, 14, 15], and visual reasoning [5].
@@
In recent years, prompt tuning [16, 17, 18, 19] has emerged as a parameter-efficient approach to fine-tune VLMs. The introduction of a few learnable prompts to a frozen pre-trained VLM (such as CLIP) allows it to adapt to various downstream tasks. However, those prompts tend to overfit the task-specific data distributions, raising concerns about their generalization capabilities across diverse tasks.
@@
To address the limited generalization capability of prompts, we consider the label smoothing (LS) [20, 21, 22], a widely recognized regularization technique. Originally introduced to mitigate the over-fitting in classification tasks, LS reduces the over-confidence of a model by distributing a small portion of probability mass across incorrect labels. LS has been shown to improve the generalization by encouraging models to produce softer and more calibrated predictions [23, 24, 25]. Given its effectiveness in reducing over-fitting and enhancing generalization, this motivates us to explore the integration of LS with prompt tuning.
@@
Figure 1: Comparison of test accuracies among CoOp [16], CoOp with vanilla label smoothing (LS) [20], TFKD [22], OLS [23], and alternating training-based label smoothing (ATLaS), where prompts are trained on the base classes of ImageNet and subsequently evaluated on new classes.
@@
However, our preliminary investigation reveals that employing the LS strategy for prompt tuning fails to enhance the generalization capability of prompts, which aligns with an observation in [26] that “label smoothing doesn’t improve fine-tuning of a pre-trained model”. Specifically, we conduct preliminary experiments using a representative prompt tuning method, CoOp [16], as a baseline, where experimental settings are described in Section IV-B3 of the base-to-new generalization part. As illustrated in Fig. 1, CoOp with vanilla LS demonstrates slightly inferior testing performance compared to the CoOp method throughout the prompt tuning process. The variants of LS do not improve the testing performance (i.e., TFKD [22]), but rather impair it (i.e., OLS [23]). Naturally, we raise a question: “Can the LS be revised to help improve the generalization capacity of prompt tuning”.
@@
To answer this question, we propose an Alternating Training-based Label Smoothing (ATLaS) method, a novel label smoothing technique, for prompt tuning. Specifically, ATLaS alternates between using one-hot labels and soft labels as the supervision information to guide prompt tuning, thereby enhancing the generalization capability of learnable prompts. Fig. 1 shows that CoOp with ATLaS successfully enhances the prompt generalization. For the generation of soft labels in VLMs, we propose two efficient methods. The first method produces Class-wise Soft Labels (CSL) to derive inter-class similarities from textual prompts, and the second method generates Instance-wise Soft Labels (ISL) to capture instance-class associations from the interaction between textual prompts and instances. Those two types of soft labels could effectively guide the learning of prompts by providing rich supervision signals from different perspectives. Theoretically, we analyze ATLaS from the optimization perspective, demonstrating its advantages over the vanilla LS. Extensive experiments demonstrate that the proposed ATLaS method effectively enhances the generalization of prompts in several settings (i.e., cross-dataset transfer, domain generalization, base-to-new, and few-shot classification).
The main contributions of this work are four-fold.
@@
  • To the best of our knowledge, this is the first attempt to successfully integrate LS into prompt tuning to improve the generalization capacity. To address the ineffectiveness of vanilla LS in prompt tuning, we propose the ATLaS method to guide prompt tuning by alternating between one-hot labels and soft labels.
  • We introduce two types of soft labels to guide prompt tuning, including class-wise and instance-wise soft labels, both of which can be efficiently generated before the training process.
  • The convergence properties of the proposed ATLaS method are analyzed.
  • Extensive experiments demonstrate the effectiveness of the proposed ATLaS method.
II Related Work
II-A CLIP-based VLMs
@@
VLMs [1, 3, 13, 5] have significantly enhanced cognitive capabilities by integrating visual and textual modalities, excelling at real-world vision understanding tasks such as image classification. Among those models, CLIP [1] stands out as a foundational pioneer, establishing the standard practice of aligning visual and textual embeddings within a unified representation space. Given CLIP’s foundational role in inspiring a broad family of subsequent models, including ALIGN [2], EVA-CLIP [4], OpenCLIP [27], and LongCLIP [28], we use CLIP as the foundational VLM to explore the novel prompt tuning paradigm with label smoothing.
II-B Prompt Tuning for VLMs
@@
As a lightweight model adaptation method, prompt tuning has attracted widespread attention in VLMs [29, 30, 31, 32]. Unlike traditional fine-tuning methods that update all parameters of the pre-trained model, prompt tuning optimizes only the prompt representations, thereby reducing the computational costs. In particular, CoOp [16] achieves efficient adaptation by optimizing a set of learnable prompt vectors, which replace the hand-crafted textual prompts (e.g., “a photo of panda”) at the language branch of CLIP. However, despite its simplicity and effectiveness, CoOp tends to overfit base classes, exhibiting limited generalization ability. To address this limitation, CoCoOp [33] learns a lightweight meta-net that generates input-conditional vectors for each input image. ProGrad [29] regulates the gradient direction to mitigate conflicts between hand-crafted prompts and learnable prompts. MaPLe [34] introduces a multi-modal prompt learning method that simultaneously learns hierarchical prompts for both vision and language branches to improve the alignment between the visual and textual representations. To address the base-new trade-off issue, DePT [35] decouples base-specific knowledge and task-shared knowledge in feature channels.
II-C Label Smoothing
@@
Empirical evidence proves that applying LS strategies during training improves the generalization of the model [20]. LS is a regularization strategy that replaces one-hot labels with soft labels, preventing the model from becoming over-confident in its predictions. Müller et al. [21] demonstrate that LS encourages samples from the same class to group in tight clusters, which results in a loss of information in the logits regarding the similarities of different classes, consequently impairing knowledge distillation. To generate more reliable soft labels, Online label smoothing (OLS) [23] derives class-wise soft labels online by retaining historical predictions. Yuan et al. [22] explore the intrinsic relationship between LS and knowledge distillation, and propose a teacher-free knowledge distillation (TFKD) framework, where a student model engages in self-learning or draws insights from manually designed regularization distribution. Different from existing LS methods whose optimization paradigm is to jointly optimize a designed loss of label smoothing along with the standard cross-entropy loss, the proposed ATLaS method alternately optimizes the model with the loss of label smoothing and the standard cross-entropy loss, achieving better performance.
III Methodology
@@
III-A Preliminaries
Contrastive Language-Image Pre-training (CLIP) [1] is a VLM pre-trained on about 400 million image-text pairs, resulting in remarkable zero-shot image recognition capability. CLIP enables zero-shot image classification by measuring the similarity between visual and textual embeddings. Specifically, CLIP consists of a visual encoder and a text encoder. Let $\mathbf{I}$ be the visual embedding extracted by the visual encoder in CLIP for an image $\mathbf{x}$, and $\{\mathbf{t}_{c}\}_{c=1}^{C}$ be a set of textual embeddings generated by the text encoder for textual prompts, where $C$ is the number of classes. Those prompts typically follow the format like “a photo of a [CLS]”, where “[CLS]” denotes a class token which is subsequently replaced by specific class names such as “panda”, “dog”, and “bird”. Given those embeddings, the prediction probability for image classification can be computed as
@@ III-A Preliminaries
Context Optimization (CoOp) [16] is a prompt tuning method designed to replace textual prompts with learnable prompts, enabling VLMs to adapt to downstream tasks through parameter-efficient fine-tuning with a few labeled samples. Specifically, CoOp introduces learnable prompt vectors $\mathbf{v}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{M}\}$ to replace the textual prompt (e.g., “a photo of a”), where each $\mathbf{v}_{i}$ has the same dimension as the word embeddings. The learnable prompt is subsequently constructed by concatenating $\mathbf{v}$ with the class token. Formally, the learnable prompts are formed as $\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{M},[\mathrm{CLS}]\}$. Let $\{\mathbf{t}_{c}^{\mathrm{s}}\}_{c=1}^{C}$ be a set of textual embeddings generated by the text encoder for the learnable prompts. The learnable prompt can be optimized by minimizing the cross-entropy loss as
@@
Label Smoothing (LS) [20] is a regularization strategy that uses soft labels, generated by the weighted averaging of a uniform distribution and hard labels, as a substitute for hard labels during the training process. This technique is designed to mitigate the model over-confidence, thereby enhancing the generalization capabilities of the model. Specifically, LS modifies the one-hot label with the soft label $\mathbf{y}^{\mathrm{LS}}$ as
@@
where $\theta\in[0,1]$ is a smoothing parameter, $\mathbf{y}$ denotes the one-hot label, and
$\hat{\mathbf{y}}\sim\mathbb{P}_{\hat{\mathbf{y}}}$ (4)
is drawn from a distribution $\mathbb{P}_{\hat{\mathbf{y}}}$. Following existing works [20, 21, 22], we consider a uniform distribution across all $C$ classes for $\hat{\mathbf{y}}$, i.e., $\hat{\mathbf{y}}={\frac{1}{C}}\mathbf{1}_{C}$, where $\mathbf{1}_{C}\in\mathbb{R}^{C}$ is the vector of all ones.
III-B Alternating Training-based Label Smoothing
The prompts learned in CoOp exhibit limited generalization capabilities [33]. As the LS strategy is specifically proposed to enhance model generalization, this naturally motivates the integration of LS into the prompt tuning framework to enhance the generalization of the resultant prompts.
However, as discussed in the introduction (i.e., Fig. 1), directly using vanilla LS fails to improve the generalization performance and even impairs it. One possible reason is that the pre-trained CLIP model is trained with one-hot labels but not soft labels, making LS integrated with prompt tuning inconsistent with such a training process and thereby weakening the generalization capacities.
To alleviate the inconsistency of labels used in the pre-training and fine-tuning phases as discussed above, the proposed ATLaS method illustrated in Fig. 2 utilizes both the one-hot labels and soft labels. Specifically, ATLaS trains with the one-hot labels and soft labels alternately in $K$ successive training epochs, where the construction of soft labels will be introduced in Section III-D. That is, in $K$ successive training epochs, the ATLaS method learns the prompts using one-hot labels $\mathbf{y}$ in the first (K1)(K-1) epochs and then trains with the soft labels $\mathbf{y}^{\mathrm{LS}}$ in the $K$-th epoch. This alternating training strategy continues throughout the training process.
III-C Analysis
@@
We present the following assumptions about smoothness and bounded variance of stochastic gradients, which are commonly used in the non-convex optimization literature [36, 37, 38].
Assumption 1
@@
As articulated in Theorem 1, a smaller $\kappa$, which is associated with the quality of $\hat{\mathbf{y}}$ employed in LS, facilitates faster convergence. It is noteworthy that the convergence bound for vanilla LS is $\frac{2F(\mathbf{v}_{0})}{\eta T}+(1-\theta+\theta\kappa)\sigma^{2}$ [39], [^2] We provide a detailed convergence analysis of label smoothing by [39] in Appendix I-D. which is unequivocally larger than that for ATLaS when $\kappa>1$. This suggests that ATLaS can converge better during the optimization of parameters where the selection of $\hat{\mathbf{y}}$ is suboptimal (i.e., characterized by a large $\kappa$).
@@
III-D Generation of Offline Soft Labels
@@
Due to inherent factors such as pose, deformation, and lighting conditions, images within the same category can exhibit substantial intra-class variance [40, 41]. Such intra-class variance prevents CSL from adequately capturing visual differences. Moreover, the utilization of knowledge in a single modality to supervise prompt tuning could potentially disrupt the vision-language alignment learned by CLIP. To address those concerns in CSL, we propose the ISL, which utilizes both the textual and visual modalities through the interaction between textual prompts and individual instances to capture instance-class associations. Specifically, we can use the prediction probabilities of CLIP as soft labels, which are instance-dependent, and learn the instance-class relations. However, this approach can not guarantee that the value for the ground-truth class in the soft label be the largest. Therefore, the soft label in ISL is rectified to be a combination of the one-hot label $\mathbf{y}$ and the prediction probabilities as
@@
where $\alpha$ is the rectification coefficient, $\mathbf{p}(\cdot|\mathbf{x})=[p(y_{1}|\mathbf{x}),p(y_{1}|\mathbf{x}),\dots,p(y_{C}|\mathbf{x})]^{\mathrm{T}}$ denotes the prediction probabilities by CLIP, $\mathbb{I}(\cdot)$ is the indicator function, and $\delta=\mathbb{I}(\mathrm{arg}\max\mathbf{p}(\cdot|\mathbf{x})\neq\mathrm{the}\;\mathrm{ground}\mbox{-}\mathrm{truth}\;\mathrm{class})$. In other words, the rectification process defined in Eq. (6) is exclusively implemented when the class corresponding to the maximum value of the soft label diverges from the ground-truth class.
@@
In summary, the whole algorithm for the proposed ATLaS method is shown in Algorithm 1.
@@
[…]
@@
III-E Analysis of the Soft Labels Generated by CSL and ISL
Here we provide a case study with detailed explanations in Fig. 3, showing the soft labels generated by CSL and ISL for two same-class images. Firstly, CSL enhances vanilla LS by incorporating inter-class similarity knowledge. While vanilla LS assigns equal value to non-target classes, CSL generates more informed soft labels (e.g., 0.96 confidence for the target class, with varying confidence for non-target classes based on their similarities). Secondly, ISL generates instance-specific soft labels to capture intra-class variance (e.g., 0.93 and 0.60 confidence for the target class of different images). For an image containing both a human and a beagle, ISL effectively captures this contextual information and assigns 0.24 confidence to the ‘boxer’ class, showing its ability to generate more semantically rich soft labels.
Figure 3: Case study of the soft labels generated by ISL and CSL.
IV Experiments
In this section, we empirically evaluate the proposed ATLaS method.
IV-A Experimental Setup
Datasets